OpenStudy (rockycatsrule):

A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct. A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

OpenStudy (smilie2015):

y – y1 = m (x – x1) m =3 , x1 = -1 , y1 = -8 y+ 8 = 3(x + 1) y+ 8 = 3x + 3 y - 3x = -5 -3x + y = -5

OpenStudy (smilie2015):

A